# VLMQ: Efficient Post-Training Quantization for Large Vision-Language Models via Hessian Augmentation

### v1

@@ 6 Experiments
Table 5: Quantization time (in GPU hours) of 2B to 32B models using a single H100 80GB GPU. VLMQ incurs negligible overhead compared to baseline methods.
Table 5 presents the quantization time cost of VLMQ compared to baseline methods. As VLMQ is formally compatible with the GPTAQ formulation (Li et al. 2025b), it can seamlessly adopt efficiency-enhancing techniques such as Cholesky decomposition and residual decomposition. Because VLMQ incorporates only token-level importance, it results in minimal structural modifications to the optimization process. The primary additional overhead arises from a single local forward and backward pass per decoding layer, which introduces negligible latency in practice.
@@ 7 Conclusion and Limitation
We introduced VLMQ, a novel importance-aware PTQ framework for VLMs. We identified the vision over-representation that hinders the direct application of existing LLM PTQ methods to VLMs. Motivated by this insight, VLMQ explicitly leverages the enhanced Hessian yielded from the importance-aware objective. VLMQ effectively enhances the performance of quantized VLMs across various vision-language benchmarks, making large VLMs more practical for deployment under ultra-low-bit settings. In terms of limitations, our evaluation primarily focuses on image-text tasks. However, we believe VLMQ can generalize to broader applications, such as video understanding and language-only tasks, which we leave for future exploration.
@@
Appendix A Theoretical Proof
Proof of Equation 7
We begin by revisiting our importance-aware Lagrangian formulation, as presented in Equation 6, given by
@@ Appendix A Theoretical Proof
The following derivation closely follows the formal framework established in GPTAQ (Li et al. 2025b), with the key distinction being the substitution of the vanilla Hessian, residual, and activation matrix by their importance-aware counterparts, as defined in Equation 7.
Proof of Theorem 5.1
$=\Delta{\mathbf{z}}{\mathbf{p}}^{(\Delta{\mathbf{z}}),\top}.$ (22)
Proof done.
Appendix B Preliminary
Figure 5: VLM architecture.
Related Works
PTQ for LLMs. Post-training quantization (PTQ) (Lv et al. 2024; Chen et al. 2024a; Tian et al. 2024; Gong et al. 2024) has become a widely adopted technique for compressing (Li and Panda 2024; Wnag et al. 2024; Huang et al. 2025a) and accelerating large models (Hu et al. 2025; Huang et al. 2025c). For LLMs, methods such as AWQ (Lin et al. 2023) and SmoothQuant (Xiao et al. 2023) address activation outliers via smooth-based transformations. Hessian-guided approaches like OBS (Frantar and Alistarh 2022), GPTQ (Frantar et al. 2022), and GPTAQ (Li et al. 2025b) offer closed-form, layer-wise quantization solutions. More recently, rotation-based techniques (Ashkboos et al. 2024) enhance quantization-friendliness by reshaping parameter distributions, leading to improved low-bit performance. In parallel, reconstruction-based methods (Huang et al. 2024, 2025b) such as OmniQuant (Shao et al. 2023), TesseraQ (Li and Panda 2024), and CBQ (Ding et al. 2023) operate orthogonally to the above and can be integrated post-quantization for further refinement.
VLM Quantization Background
VLM architecture. As shown in Figure 5, the VLMs comprise three key components: a visual encoder, a vision-text projector, and a Language Model (LM) backbone. The visual encoders can be a pre-trained image encoder like a CLIP-style encoder (Radford et al. 2021), which is responsible for converting image/video input into vision tokens. Then they are fed into the MLP-based projector to align it with the LM embedding space. Together with encoded text tokens, they are further fed into the LM backbone. The LM backbone usually comprises multiple stacked decoding layers followed by an LM prediction head. The modality gap across hidden states in the word embedding space can be observed by PCA.
Given the multi-modal input ${\mathbf{X}}\in{\mathbb{R}}^{C_{i}\times N}$ of language model backbone, we can disassemble it based on different functionalities as
${\mathbf{X}}={\mathbf{X}}_{t}\oplus{\mathbf{X}}_{v},$ (24)
where text and vision token set ${\mathbf{X}}_{t}$ and ${\mathbf{X}}_{v}$ are
${\mathbf{X}}_{t}={\mathcal{S}}\{{\mathbf{X}}_{\texttt{sys}},{\mathbf{X}}_{\texttt{ins}},{\mathbf{X}}_{\texttt{ans}}\}$ (25)
${\mathbf{X}}_{v}={\mathcal{S}}\{{\mathbf{X}}_{\texttt{img}}\}$ (26)
Therefore, the Hessian information matrix can be written as
${\mathbf{H}}=\sum_{i=0}^{N_{t}-1}{\mathbf{x}}_{i}{\mathbf{x}}_{i}^{\top}+\sum_{j=0}^{N_{v}-1}{\mathbf{x}}_{j}{\mathbf{x}}_{j}^{\top}={\mathbf{H}}_{t}+{\mathbf{H}}_{v}$ (27)
where ${\mathbf{H}}_{t}$ and ${\mathbf{H}}_{v}$ represent the Hessian matrix formulated by text and vision tokens, respectively. Generally, the vision token count is much more than the text token count (i.e. $N_{v}\gg N_{t}$) and always exhibits high-degree of redundancy, leading to the overall feature implied by $\mathbf{H}$} biased towards vision modality.
Quantization basic. This paper focuses on uniform affine quantization. Quantization essentially maps high-precision (e.g., FP16/BF16) weights/activations to low-precision (e.g., INT8/INT4) formats to reduce memory footprint and achieve inference acceleration, which is given by
Hessian-based Quantization
OBQ calibrates the quantization of linear layers in models by iteratively 1) quantizing entries that introduce minimal loss, and 2) updating the remaining unquantized ones with the guidance of Hessian ${\mathbf{H}}={\mathbf{X}}{\mathbf{X}}^{\top}$. GPTQ (Frantar et al. 2022) extends this paradigm by formulating fixed-order entry quantization in a parallel manner, lazy block updating and Cholesky reformulation to achieve efficient calibration. GPTAQ (Li et al. 2025b) further identifies the critical error accumulation issue and proposes an asymmetric calibration principle to enhance the performance. Other Hessian-based variants (Kim et al. 2024; Edalati et al. 2025) turn to find a more accurate Hessian representation to preserve the quantized model capabilities.
@@ Appendix B Preliminary
Appendix C VLMQ
Appendix D Experiments
Implementation Details
@@ Appendix D Experiments
Calibration. We randomly sample 512 data samples as the calibration set. Each sample is structured according to Equation 23 and concatenated with the default prompt template. The input sequence length is adjusted based on the target model family: for Qwen2-VL and Qwen2.5-VL series, we truncate inputs to 512 tokens, while for LLaVA-OneVision series, the length is set to 986 tokens. To ensure semantic completeness, we discard samples whose vision token segments are truncated mid-way, as such instances result in incomplete symbolism. For quantization parameters, we adopt most settings from GPTAQ (Li et al. 2025b). Specifically, we enable act_order to enhance numerical stability. For group-wise quantization, we disable static_group to allow dynamic group-wise quantization parameter assignment.
@@
Evaluation. We implement VLMQ using the Huggingface Transformers library (Wolf et al. 2019) on top of the PyTorch framework. For evaluation, we leverage the open-source LMMs-Eval toolkit (Zhang et al. 2024a), with minor adaptations tailored to our quantized settings. During evaluation, we observe that the default answer matching mechanism in LMMs-Eval is suboptimal for ScienceQA, especially under ultra-low-bit quantization scenarios. In particular, although the model often provides the correct answer, it may not strictly adhere to the required output format, namely, returning only the option letter. For instance, a response such as “The answer is B.” semantically matches the correct answer but fails to conform to the expected literal format (e.g., “B” or “B.”), leading to a false negative in accuracy computation. To mitigate this mismatch, we introduce a robust post-processing pipeline for automatic answer normalization and extraction. This process involves parsing the model’s textual response and identifying the most probable option letter based on regular expressions and contextual cues. The implementation details of this logic are provided in Listing 1. This enhancement ensures fairer evaluation and better reflects the true capability of quantized models on instruction-following tasks such as ScienceQA.
@@
VLMQ 8.04 16.40 20.42 15.08 29.20 7.90 3.03 12.77 14.86
Table 6: INT3 performance comparison on 0.5B/32B models.
Additional results. Table 6 presents additional results on 0.5B and 32B models. It is evident that VLMQ consistently outperforms baseline approaches. However, despite its effectiveness, the 0.5B model still exhibits a notable performance gap compared to its full-precision counterpart, highlighting the challenge of quantizing smaller VLMs under ultra-low-bit settings.
Listing 1 Additional string matching logic for ScienceQA
1# pred is the model response string
2if len(pred) >= 2 and pred[1] == ".":
3 result = 1.0 if pred[0] == target else 0.0
4 return {"exact_match": result}
5if ’A.’ in pred or ’A\"’ in pred or ’A in pred or pred[-1] == ’A’:
6 result = 1.0 if target == ’A’ else 0.0
7 return {"exact_match": result}
8if ’B.’ in pred or ’B\"’ in pred or ’B in pred or pred[-1] == ’B’:
9 result = 1.0 if target == ’B’ else 0.0
10 return {"exact_match": result}
11if ’C.’ in pred or ’C\"’ in pred or ’C in pred or pred[-1] == ’C’:
12 result = 1.0 if target == ’C’ else 0.0
13 return {"exact_match": result}
14if ’D.’ in pred or ’D\"’ in pred or ’D in pred or pred[-1] == ’D’:
15 result = 1.0 if target == ’D’ else 0.0
16 return {"exact_match": result}

### v2

@@ 5.4 Efficiency
[…]
Table 10 reports the quantization latency and peak memory consumption of VLMQ, GPTQ, and GPTAQ across different model scales. Since VLMQ preserves the original GPTQ-format quantization scheme, it remains compatible with existing efficiency optimizations and introduces only token-level importance estimation as additional overhead. Consequently, the extra quantization cost is small, with latency increasing by less than 10 minutes depending on model size. The memory overhead mainly comes from storing activation gradients for one local backward pass per block, which remains modest relative to the total memory budget and well within the capacity of a single H100 80GB GPU even for 32B-scale models. Overall, the results show that VLMQ achieves its performance gains with only lightweight memory overhead and negligible impact on quantization time.
5.4.2 Inference speedup.
[…]
Linear Size of $\mathbf{W}^{\ast}$ FP16 W3 Speedup
q/k/v_proj $(3584,4608)$ 38.97 4.17 $9.34\times$
out_proj $(3584,3584)$ 39.00 4.17 $9.34\times$
gate/up_proj $(3584,37888)$ 175.17 18.46 $9.49\times$
down_proj $(18944,3584)$ 96.39 10.47 $9.20\times$
VLMQ is fully compatible with the standard GPTQ inference pipeline and can be deployed directly on existing GPTQ-compatible backends with hardware-optimized low-bit kernels [frantar2022gptq]. This compatibility is further evidenced by the latency results in Table 11. Using the official GPTQ W3 kernel on an RTX 5090 GPU, the representative linear layers in Qwen2-VL-7B-Instruct exhibit substantial speedups over FP16 execution. These results confirm that VLMQ preserves the same inference efficiency as GPTQ and can seamlessly inherit the benefits of existing optimized low-bit deployment stacks, without introducing any additional inference overhead.
@@ 6 Conclusion and Limitation
We introduced VLMQ, a novel importance-aware PTQ framework for VLMs. We identified the vision over-representation that hinders the direct application of existing LLM PTQ methods to VLMs. Motivated by this insight, VLMQ explicitly leverages the enhanced Hessian yielded from the importance-aware objective. VLMQ effectively enhances the performance of quantized VLMs across various vision-language benchmarks, making large VLMs more practical for deployment under ultra-low-bit settings. In terms of limitations, our evaluation primarily focuses on image-text tasks. However, we believe VLMQ can generalize to broader applications, such as video understanding tasks, which we leave for future exploration.
@@
Appendix Overview
We provide additional details in the appendix, organized as follows
  • $\triangleright$
    Section 0.A supplements related works on Hessian-based quantization for LLMs. In particular, we summarize the motivation, methodology, and limitations of existing approaches, and provide an overview of GPTAQ [li2025gptqv2] as the baseline precursor algorithm adopted in this work.
  • $\triangleright$
    Section 0.B describes the detailed architectures of representative VLMs and outlines the necessary quantization preliminaries.
  • $\triangleright$
    Section 0.C elaborates on the proposed VLMQ framework. We provide a step-by-step derivation of the optimization objective and the role of importance factors. For reproducibility, we also include detailed pseudo-code illustrating the complete quantization pipeline.
  • $\triangleright$
    Section 0.D presents the theoretical justification of our method, including the full proof of the theorem introduced in the main paper.
  • $\triangleright$
    Section 0.E reports extended experimental settings and comprehensive results. This includes dataset descriptions, evaluation metrics, implementation details, and additional evaluation results that complement the main results and further validate the effectiveness of our approach.
Appendix 0.A Related Works
0.A.0.1 Hessian-based LLM quantization.
Hessian information is widely used as a calibration guide to compensate for quantization errors. OBQ [frantar2022optimal-brain-compression] calibrates the quantization of linear layers in models by iteratively 1) quantizing entries that introduce minimal loss, and 2) updating the remaining unquantized ones with the guidance of Hessian ${\mathbf{H}}={\mathbf{X}}{\mathbf{X}}^{\top}$. GPTQ [frantar2022gptq] extends this paradigm by formulating fixed-order entry quantization in a parallel manner, lazy block updating, and Cholesky reformulation to achieve efficient calibration. GPTAQ [li2025gptqv2] further extends GPTQ by optimizing an asymmetric objective:
[…]
where $q$ denotes the index of the quantized weight element. GPTAQ thereby achieves accurate quantization while better preserving model functionality, with efficiency ensured by techniques such as efficient residual decomposition. Other works [kim2024boa, edalati2025oac, guan2024aptq] focus on obtaining accurate Hessian information to enhance PTQ, enable mixed-precision quantization, and support related techniques.
Appendix 0.B VLM Quantization Background
0.B.1 VLM Architecture
As shown in Figure 6, the VLMs comprise three key components: a visual encoder, a vision-text projector, and a Language Model (LM) backbone. The visual encoders can be a pre-trained image encoder like a CLIP-style encoder [radford2021clip], which is responsible for converting image/video input into vision tokens. Then they are fed into the MLP-based projector to align it with the LM embedding space. Together with encoded text tokens, they are further fed into the LM backbone. The LM backbone usually comprises multiple stacked decoding layers followed by an LM prediction head. The modality gap across hidden states in the word embedding space can be observed by PCA.
Figure 6: VLM architecture.
Given the multi-modal input ${\mathbf{X}}\in{\mathbb{R}}^{C_{i}\times N}$ of the language model backbone, we can disassemble it based on different functionalities as
${\mathbf{X}}={\mathbf{X}}_{t}\oplus{\mathbf{X}}_{v},$ (10)
where text and vision token set ${\mathbf{X}}_{t}$ and ${\mathbf{X}}_{v}$ are ${\mathbf{X}}_{t}={\mathcal{S}}\{{\mathbf{X}}_{\texttt{sys}},{\mathbf{X}}_{\texttt{ins}},{\mathbf{X}}_{\texttt{ans}}\}$ and ${\mathbf{X}}_{v}={\mathcal{S}}\{{\mathbf{X}}_{\texttt{img}}\}$, respectively.
0.B.2 Quantization Preliminary
This paper focuses on uniform affine quantization. Quantization essentially maps high-precision (e.g., FP16/BF16) weights/activations to low-precision (e.g., INT8/INT4) formats to reduce memory footprint and achieve inference acceleration, which is given by
Appendix 0.C VLMQ Framework
0.C.1 Refined Objective
In this paper, we choose the advanced GPTAQ as our precursor algorithm. Although equation 8 yields the optimal quantized weight $\hat{{\mathbf{w}}}$ by minimizing the squared error, it inherently assumes equal contribution across all tokens in ${\mathbf{Z}}={\mathbf{W}}{\mathbf{X}}$. Such equal treatment may introduce bias due to the vision over-representation. To address this limitation, we introduce a token-level importance weighting matrix ${\mathbf{G}}\in{\mathbb{R}}^{N\times N}$, where $\mathbf{G}$} is diagonal with the $i$-th diagonal element representing the importance assigned to output token ${\mathbf{Z}}_{:,i}$. Incorporating importance factors into the objective leads to a refined formulation:
[…]
@@ 0.C.1 Refined Objective
The following derivation closely follows the formal framework established in GPTAQ [li2025gptqv2], with the key distinction being the substitution of the vanilla Hessian, residual, and activation matrix by their importance-aware counterparts, as defined in equation 19. We present the corresponding optimal weight updating in VLMQ as
$\Delta{\mathbf{w}}$ $=\frac{\left(\hat{{\mathbf{w}}}_{q}-{\mathbf{w}}_{q}\right)}{\left[{\mathbf{X}}{\mathbf{G}}^{2}{\mathbf{X}}^{\top}\right]_{qq}^{-1}}\cdot\widetilde{{\mathbf{H}}}_{q:,}^{-1}+{\mathbf{r}}{\mathbf{G}}^{2}{\mathbf{X}}^{\top}\left[{\mathbf{X}}{\mathbf{G}}^{2}{\mathbf{X}}^{\top}\right]_{-q}^{-1}$ (19)
where $\widetilde{{\mathbf{H}}}={\mathbf{X}}{\mathbf{G}}^{2}{\mathbf{X}}^{\top}$, $\tilde{{\mathbf{r}}}={\mathbf{r}}{\mathbf{G}}$, and $\widetilde{{\mathbf{X}}}={\mathbf{X}}{\mathbf{G}}$ indicate the importance-aware Hessian, residual, and activation matrix yielded from the refined objective (equation 15). Our importance-aware weight update presented in equation 19 remains formally consistent with the original GPTAQ [li2025gptqv2] formulation. Distinct from BoA [kim2024boa], which attributes importance at the channel level, our approach assigns importance scores at the token-level activation granularity. This design is orthogonal to the parallelization strategies proposed in GPTQ [frantar2022gptq] and GPTAQ [li2025gptqv2]. The compatibility allows us to directly inherit efficiency tricks (e.g. Cholesky reformulation) proposed therein.
0.C.2 VLMQ
@@
Appendix 0.D Theoretical Proof
0.D.1 Proof of Theorem 4.1
$=\Delta{\mathbf{z}}{\mathbf{p}}^{(\Delta{\mathbf{z}}),\top}.$ (27)
Proof done.
Appendix 0.E Experiments
0.E.1 Implementation Details
0.E.1.1 Calibration.
@@ 0.E.1 Implementation Details
We randomly sample 512 data samples from COCO Caption datasets [chen2024sharegpt4v] as the calibration set. Each sample is structured according to equation 9 and concatenated with the default prompt template. The input sequence length is adjusted based on the target model family: for Qwen2-VL and Qwen2.5-VL series, we truncate inputs to 512 tokens, while for LLaVA-OneVision series, the length is set to 986 tokens. To ensure semantic completeness, we discard samples whose vision token segments are truncated mid-way, as such instances result in incomplete symbolism. For quantization configurations, we adopt most settings from GPTAQ [li2025gptqv2]. Specifically, we enable act_order to enhance numerical stability. For group-wise quantization, we disable static_group to allow dynamic group-wise quantization parameter assignment.
0.E.1.2 Evaluation.
@@
We implement VLMQ using the Huggingface Transformers library [wolf2019huggingface] on top of the PyTorch framework. For evaluation, we leverage the open-source LMMs-Eval toolkit [zhang2024lmms-eval], with minor adaptations tailored to our quantized settings. During evaluation, we observe that the default answer matching mechanism in LMMs-Eval is suboptimal for ScienceQA, especially under ultra-low-bit quantization scenarios. In particular, although the model often provides the correct answer, it may not strictly adhere to the required output format, namely, returning only the option letter. For instance, a response such as “The answer is B.” semantically matches the correct answer but fails to conform to the expected literal format (e.g., “B” or “B.”), leading to a false negative in accuracy computation. To mitigate this mismatch, we introduce a robust post-processing pipeline for automatic answer normalization and extraction. This process involves parsing the model’s textual response and identifying the most probable option letter based on regular expressions and contextual cues.
0.E.2 Additional results.
0.E.2.1 Additional illustration on INT2g128 quantization.
While taking GPTAQ as our precursor algorithm generally yields satisfactory results under INT3g128 quantization, applying this configuration to INT2g128 leads to severe performance degradation. In particular, GPTAQ produces even worse results than GPTQ under this ultra-low-bit setting. Taking Qwen2.5-VL-7B-Instruct-INT2g128 as an example, GPTAQ achieves only 41.69% accuracy, significantly lower than the 55.58% achieved by GPTQ. Additionally, the Qwen2.5-VL-7B-Instruct-INT2g128 fails on MME-RealWorld (Chinese) and ScienceQA benchmarks, achieving only 3.60% and 2.95% accuracy, respectively. A plausible explanation is that, under extremely low-bit quantization, the residual error in GPTAQ propagated from previously quantized layers may be continuously accumulated and significantly amplified. This accumulated noise severely interferes with the dominant weight update, thereby compromising the undesirable quantization quality. Therefore, in cases where GPTAQ underperforms GPTQ, we opt to use GPTQ as our precursor algorithm. It is important to note that, since our proposed VLMQ framework is designed as a plug-and-play solution for VLMs, it allows for seamless adaptation from a GPTAQ-based implementation to one based on GPTQ.
Table 12: INT3 quantization performance comparison across different quantization methods on eight benchmarks. indicates that we use the token-level $\ell_{2}$-norm of row gradients instead of $\ell_{1}$-norm as our importance factors.
[…]
Qwen2-VL-2B-Instruct-INT3
Full Prec 72.61 89.35 40.27 43.59 76.60 74.23 61.79 79.38 67.23
AWQ 5.36 10.02 9.65 8.48 5.40 1.37 18.53 4.37 7.90
MBQ 3.56 10.19 10.22 7.79 4.40 1.44 8.12 3.90 6.20
GPTQ 58.84 74.55 33.25 30.83 61.50 65.13 50.37 73.49 56.00
GPTAQ 58.76 74.03 38.71 28.71 61.10 64.25 52.17 73.18 56.36
[…]
Qwen2-VL-7B-Instruct-INT3
Full Prec 81.44 93.89 57.30 56.09 80.90 85.55 69.26 82.02 75.81
AWQ 18.20 27.67 30.20 20.03 8.90 49.89 44.88 28.23 28.50
MBQ 16.52 24.75 27.23 21.35 8.10 35.25 40.10 27.79 25.13
GPTQ 72.80 88.35 47.76 27.90 72.40 78.26 63.02 78.28 66.10
GPTAQ 72.76 88.09 49.76 43.96 73.80 80.41 64.21 79.23 69.03
[…]
LLaVA-OneVision-7B-INT3
Full Prec 80.08 87.09 57.39 53.93 62.10 89.98 64.82 75.96 71.42
GPTQ 74.24 77.82 46.47 42.59 56.60 85.26 60.65 71.90 64.44
GPTAQ 75.72 78.07 46.48 42.94 58.30 85.97 61.84 72.69 65.25
[…]
Qwen2.5-VL-7B-Instruct-INT3
Full Prec 83.20 94.72 58.55 52.98 84.40 88.26 70.62 83.05 76.97
GPTQ 62.44 89.33 42.31 27.33 77.70 79.39 66.58 78.38 65.43
GPTAQ 63.68 90.41 47.18 36.49 79.60 80.52 66.40 77.94 67.78
@@
[…]
Table 13: INT4 quantization performance comparison across different quantization methods on eight benchmarks.
[…]
Qwen2-VL-7B-Instruct-INT4
GPTQ 80.44 93.16 55.44 44.03 79.70 84.08 67.63 81.55 73.25
GPTAQ 79.88 93.04 55.55 47.30 79.50 84.23 67.85 81.68 73.63
VLMQ 80.28 93.23 55.88 47.73 80.30 84.30 67.72 81.48 73.87
0.E.2.2 Results of INT3/INT4 quantization.
We assess the zero-shot performance of INT3/INT4 quantized models across eight vision-language benchmarks in Table 12 and Table 13. For INT3 quantization, VLMQ demonstrates overall superiority compared to baseline approaches and achieves up to 1.01% average accuracy improvement. For Qwen2.5-VL-7B-Instruct-INT3, the proposed VLMQ outperforms its counterparts quantized by GPTQ and GPTAQ on all benchmarks except MME-RealWorld (English). Remarkably, it yields accuracy improvements of 1.64%, 1.40%, and 1.18% on ChartQA, MME-RealWorld (Chinese), and ScienceQA, respectively. For the LLaVA-OneVision-7B-INT3 model, we apply the token-level $\ell_{2}$-norm instead of the $\ell_{1}$-norm adopted in the other experiments, which we empirically find beneficial for performance enhancement. However, despite its effectiveness, the 0.5B model still exhibits a notable performance gap compared to its full-precision counterpart, highlighting the challenge of quantizing smaller VLMs under ultra-low-bit settings. We leave it as our future work. For the INT4 quantized model, VLMQ also delivers solid performance, showing small but consistent improvements over existing baselines. This confirms that the method remains effective even under moderate-bit quantization.
0.E.2.3 Extended ablation studies.
We provide the detailed results of Table 6, Table 7, and Table 8 in Table 14, Table 15, and Table 16, respectively. The label corresponds to the line number in the compact table.
Table 14: Ablation study on the importance factor type. Reported results are under the INT-3 setting.
[…]
Qwen2-VL-7B-Instruct, Avg Acc: 75.82%
72.76 88.09 49.76 43.96 73.80 80.41 64.21 79.23 69.03
69.40 88.32 49.88 45.28 72.90 78.54 63.64 78.33 68.29
71.60 86.83 48.37 40.76 70.20 78.90 63.20 76.71 67.07
72.76 88.90 50.17 47.25 73.50 80.05 65.52 79.41 69.70
Table 15: Ablation study on backward granularity. Reported results are under the INT-3 setting.
[…]
Qwen2-VL-7B-Instruct, Avg Acc: 75.82%
71.08 87.18 48.23 38.57 73.70 79.81 64.43 79.14 67.77
69.80 87.72 49.35 45.17 73.70 80.62 64.21 78.54 68.64
72.76 88.90 50.17 47.25 73.50 80.05 65.52 79.41 69.70
Table 16: Ablation study on layers to enhance under the INT-3 setting.
[…]
Qwen2-VL-7B-Instruct, Avg Acc: 75.82%
72.76 88.09 49.76 43.96 73.80 80.41 64.21 79.23 69.03
58.16 73.21 33.51 29.15 62.80 61.71 52.88 72.88 55.54
53.68 74.22 37.96 35.42 61.80 62.53 50.42 73.62 56.21
72.76 88.90 50.17 47.25 73.50 80.05 65.52 79.41 69.70
0.E.3 Compatibility With Hardware-Optimized Kernels
Regarding deployment efficiency on hardware, we stress that VLMQ is fully compatible with GPTQ’s quantization format. As a result, it can be seamlessly incorporated into existing works and directly utilize all hardware-optimized kernels and infrastructures designed for GPTQ, without introducing extra inference overhead. This compatibility further allows VLMQ to immediately take advantage of specialized kernels such as Marlin [frantar2024marlin] and ExLLaMA [exllamav2], eliminating the need for additional hardware optimizations.